OpenStudy (sleepyjess):

A projectile is thrown upward so that its distance above the ground after t seconds is given by the function h(t) = - 16t^2 + 640t. After how many seconds does the projectile take to reach its maximum height? Show your work.

OpenStudy (sleepyjess):

@phi

OpenStudy (sleepyjess):

@Hero @mathstudent55

OpenStudy (phi):

h(t) = - 16t^2 + 640t if you write it as \[ y = - 16x^2 + 640x \] to make it look more familiar (?), that is the equation of a parabola, with a "frown" shape \( \cap \) Its "peak" is at its vertex. Do you remember how to find the x value of the vertex? it is -b/(2a) when the equation is in the form \[ y = a x^2 +bx + c\]

OpenStudy (sleepyjess):

-640/-32?

OpenStudy (phi):

yes, which is ?

OpenStudy (sleepyjess):

20

OpenStudy (phi):

yes. x=20. Or, going back to the original equation \[ h(t) = - 16t^2 + 640t \] t=20 seconds that is how many seconds the projectile takes to reach its maximum height

OpenStudy (phi):

they don't ask, but the height would be h(20)= -16*20*20 + 640*20 = 6400

OpenStudy (sleepyjess):

Thank you so much!

OpenStudy (phi):

yw